OpenStudy (anonymous):

Solve the equation 4x+8/2 = -10 for x.

OpenStudy (lilg132):

you solve it and i'll help you

OpenStudy (lilg132):

i like to work with numbers instead of fractions so 8/2 = 4 4x+4 = -10 does that make it easier?

OpenStudy (anonymous):

2x+4=-10 2x=-14 x=-7?

OpenStudy (lilg132):

how did you get 2x?

OpenStudy (anonymous):

I divided 4x by 2

OpenStudy (lilg132):

why where did you get 2 from?

OpenStudy (lilg132):

are you still there

OpenStudy (anonymous):

Yeah

OpenStudy (lilg132):

where did you get 2 from? i just want to know so i can help you with where u are making a mistake

OpenStudy (lilg132):

you got the steps right you just made that one mistake which im trying to find out where you got 2 from?

OpenStudy (anonymous):

The equation is 4x+8 over 2 = -10

OpenStudy (lilg132):

i replaced 8/2 with 4 8/2 is the same as 4

OpenStudy (anonymous):

im confused

OpenStudy (lilg132):

ok lets start again

OpenStudy (lilg132):

what is 8/2 as a number

OpenStudy (lilg132):

what is 8 divided by 2

OpenStudy (lilg132):

u still there?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

Solve the equation 4x+8/2 = -10 for x hint : multiply whole equation by 2 and solve for x. Will make much easier for you to see.